Your Company is considering a new project that will require $540,000 of new equipment at the start of the project. The equipmI am unsure what the answer is for this question.
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(540,000-100,000)/10

=$44,000

Hence annual tax benefits=Annual depreciation*Tax rate

=44000*21%

=$9240

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=9240[1-(1.13)^-10]/0.13

=9240*5.42624348

=$50138(Approx).